Find an equation of the line having the specified slope and containing the indicated point. Write your answer in​ slope-intercept form.
mequals=six sevenths
6
7​; left parenthesis 0 comma negative 8 right parenthesis

Found 2 solutions by Alan3354, rothauserc:
Answer by Alan3354(69443)

Use math symbols.

Answer by rothauserc(4718)

m = (6/7) and given point (0, -8)
:
slope intercept form is
:
y = mx + b, where m is the slope an b is the y intercept
:
using given point substitute for x and y
:
-8 = (6/7)*0 + b
:
b = -8
:
************
y = 6x/7 - 8
************
: